Given that 𝑓(π‘₯) = 3π‘₯ 2 βˆ’ 4π‘₯ + 7, use the definition of the derivative to find 𝑓 β€² (π‘₯)


1
Expert's answer
2021-06-07T11:51:10-0400

Solution:

"\ud835\udc53(\ud835\udc65) = 3\ud835\udc65^2 \u2212 4\ud835\udc65 + 7"

By definition of derivatives:

"\begin{aligned}
f^{\prime}(x) &=\lim _{h \rightarrow 0} \frac{f(x+h)-f(x)}{h} \\
&=\lim _{h \rightarrow 0} \frac{\left(3(x+h)^{2}-4(x+h)+7\right)-\left(3 x^{2}-4 x-7\right)}{h} \\
&=\lim _{h \rightarrow 0} \frac{\left(3\left(x^{2}+2 x h+h^{2}\right)-4 x-4 h\right)-\left(3 x^{2}-4 x\right)}{h} \\
&=\lim _{h \rightarrow 0} \frac{3 x^{2}+6 x h+3 h^{2}-4 x-4 h-3 x^{2}+4 x}{h} \\
&=\lim _{h \rightarrow 0} \frac{6 x h+3 h^{2}-4 h}{h} \\
&=\lim _{h \rightarrow 0}(6 x+3 h-4) \\
&=6 x-4
\end{aligned}"